OpenStudy (anonymous):

1. what is the value of n so that the expression x^2+12x+n is a perfect square trinomial a.6 b.36 c.72 d.144 2. what are the solutions of the equation x^2+4x=96 a.2,-48 b.-12,8 c.-2,48 d12,-8

OpenStudy (loser66):

Any idea?

OpenStudy (anonymous):

I think 1 is b and I have no idea on 2 @Loser66

OpenStudy (luigi0210):

Yes, you would be right on 1, to get it to be a perfect square you take b and.. \(\Large (\frac{ b}{2})^2 \) For 2, set it equal to zero to solve.

OpenStudy (anonymous):

so it would be c

OpenStudy (luigi0210):

c? Nope: x=-2, 48 \((x+2)(x-48)=0 \) \(x^2-48x+2x-96\) \(\Large =x^2-46x-96\))